A boat went 10km upstream against the current then turned around and went 15km downstream with current. The current speed was 3km/hour and upstream trip took 30 minutes longer. What was the speed of the boat without the speed of the current?
Answer by mananth(16949)

10km upstream against the current
15km downstream with current.
current speed was 3km/hour
upstream trip took 30 minutes longer.
..
let speed in still water be x km/hour
..
downstream speed = x+3 km/hour
15 km distance.
time taken = 15/ x+3
..
upstream speed will be x-3 km/hour
distance 10 km.
time taken = 10/x-3
..
upstream time - downstream time = 1/2 hour
10/ x-3 - 15/ x+3 =1/2
LCM of left side
(x+3)(x-3)
10(x+3)-15(x-3)/ (x+3)(x-3)= 1/2
10x+30-15x+45=1/2 * (x+3)(x-3)
-5x+75= 1/2 * x^2-9
2(-5x+75)=x^2-9
-10x+150=x^2-9
x^2+10x-159=0
..
the roots of the equation x1 , x2 will be
x1= (-10+sqrt(100+636) )/2
x1=-18.56 since it is negative ignore.
x2=(-10-sqrt(100+636) )/2
x2= 8.56 km/hr the speed of boat in still water